OpenStudy (anonymous):

Integrate from 0 to 1 sin(pix)dx Answer: 2/pi Please show the work on how to get the answer.

OpenStudy (anonymous):

try giving u substitution a shot

OpenStudy (anonymous):

should i make my u pix or just x and bring pi out to the front?

OpenStudy (tkhunny):

Can you do \(\int \sin(x)\;dx\)? How about \(\int \sin(2x)\;dx\)?

OpenStudy (anonymous):

i integrated it and got pi(-cosx) from 0 to 1. how would i find cos(1) without using a calculator?

OpenStudy (anonymous):

so if you choose your u to be pi*x, then du = pi*dx therefore du/pi = dx then evaluate the integral with u, then either change your bounds or re substitute in pi*x for your u

OpenStudy (anonymous):

oh that's what i forgot! i didn't put my coordinates in terms of u x0 thanks!

OpenStudy (anonymous):

no worries!